# Discrete Diffusion in Large Language and Multimodal Models: A Survey

### v3

@@ IV.E Addressing Training-Testing Discrepancy
IV.F Reweighting Technique
IV.F.1 Multi-Granularity Diffusion Modeling
@@
IV.G Distillation through Dimensional Correlations
@@ IV.G.1 Distillation Loss
$\mathcal{L}_{\text{distil}}(\theta;\psi,r_{\delta},\delta)$ $=\mathbb{E}_{x_{\delta}\sim r_{\delta}}\Big{[}\mathrm{D_{KL}}\big{(}p^{\psi}_{0|\delta}(\cdot\mid x_{\delta})$
$\qquad\|\,p^{\theta}_{0|\delta}(\cdot\mid x_{\delta})\big{)}\Big{]}$ (80)
@@ IV.G.2 Consistency Loss
$\mathcal{L}_{\text{consis}}(\theta;\psi,r_{t},s,u,t)={}$ $\mathbb{E}_{x_{t}\sim r_{t}}\Big{[}\mathrm{D_{KL}}\bigl{(}(p^{\theta}_{s|u}\circ p^{\psi}_{u|t})(\cdot\mid x_{t})$
$\|\,p^{\theta}_{s|t}(\cdot\mid x_{t})\bigr{)}\Big{]}$ (81)
@@
V Inference Techniques
V.A Unmasking Techniques
TABLE I: The inference techniques used during the inference process of each model.
Inference Technique Language Model Multimodal Model Unified Model
LLaDA [6] DIFFUSION-LLMs [42] DiffuLLaMA [18] DREAM [7] LLaDA 1.5 [43] TESS 2 [44] Dimple [8] LaViDa [10] LLaDA-V [9] MMaDA [11] FUDOKI [45] Muddit [46]
Unmasking Top-$s_{t}$ Strategy
Confident Decoding
Semi-Autoregressive Decoding
Prefilling
Reward Model Guidance
@@ V.A Unmasking Techniques
In dLLMs and dMLLMs, the model predicts all the response tokens at each step. However, only a subset of the masked tokens are selected to be unmasked at every iteration, while the remainder remain masked. This iterative unmasking process gradually reveals the complete output sequence. The core challenges in this procedure are: (1) determining which tokens to unmask at each iteration, and (2) deciding how many tokens should be unmasked at each iteration. Table I provides a summary of the unmasking strategies adopted by representative models. Figure 5 provides a detailed illustration of each type of unmasking strategy. In this section, we discuss each category in detail and describe the specific unmasking strategies proposed in each work.
V.A.1 Discrete-Time Unmasking
@@
Let $p\in\mathbb{R}^{K}$ be the predicted probability distribution over the vocabulary for a given token, where $K$ is the vocabulary size. The following metrics are commonly used [7]:
  • Maximum Probability (Confidence):
    $c=\max(p),$ (82)
@@
  • Margin:
    $c=p_{\text{top1}}-p_{\text{top2}},$ (83)
    where $p_{\text{top1}}$ and $p_{\text{top2}}$ are the first and second highest probabilities, respectively. This measures how dominant the top prediction is.
  • Negative Entropy:
    $c=-\sum_{i=1}^{K}p_{i}\log(p_{i}+\epsilon),$ (84)
@@
  • Local Unmasking: The above strategies can generally be categorized as global unmasking strategies. When selection requires sorting, global unmasking strategies aggregate the metric values of all tokens and perform sorting or selection on this global set. However, the unmasking process does not necessarily have to be global; it can also be local. Specifically, all tokens can be divided into multiple subgroups, with sorting and selection performed independently within each subgroup, or different unmasking strategies applied across different subgroups. A typical example is the semi-autoregressive decoding strategy [6]. This approach divides the full response into multiple blocks, similar to block diffusion. During each forward pass, predictions are generated for all blocks simultaneously. However, the unmasking of tokens follows a left-to-right, block-by-block order. Tokens in the next block are only allowed to be unmasked once all tokens in the previous block have been unmasked.
@@
V.A.2 Continuous-Time Unmasking (Flow Matching)
@@
V.B Remasking Techniques
For discrete diffusion models based on absorbing states, during inference, once a token is unmasked (i.e., decoded), it remains unchanged in subsequent steps. This static behavior limits the model’s capacity to revise or refine earlier predictions. To address this, the remasking technique reintroduces masked tokens at previously unmasked positions, enabling iterative refinement of generated outputs. Remasking is able to extend the number of decoding steps beyond the response length, thereby allowing repeated updates to the response. This mechanism serves as a form of test-time scaling, improving output quality through progressive correction.
V.B.1 Remasking in General Masked Diffusion Models.
@@ V.B Remasking Techniques
V.B.2 Remasking under Discrete Flow Matching.
@@
V.C Prefilling and Caching Techniques
@@
Prefilling and Key-Value Cache (KV-Cache) are standard inference acceleration techniques widely adopted in autoregressive language models. Intuitively, Prefilling and KV-Cache avoids redundant computation by storing the key and value representations from previous decoding steps, enabling the model to reuse them instead of recalculating at each new time step.
Formally, in a transformer decoder, the current hidden states $\mathbf{h}_{i}$ are projected into a query-key-value triplet $(\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i})$ using learned projection matrices. For the first generation iteration $i=1$, the hidden state $\mathbf{h}_{0}$ corresponding to the first predicted token and all prompt tokens is used to compute the new query-key-value triplet $(\mathbf{Q}_{1},\mathbf{K}_{1},\mathbf{V}_{1})$. The calculated key tensor $\mathbf{K}_{1}$ and value tensor $\mathbf{V}_{1}$ are cached for reusing. At iteration $i>1$, only the hidden state $\mathbf{h}_{i}$ corresponding to the $i$-th response token is used to compute the new query-key-value triplet. The new key $\mathbf{K}_{i}$ and new value $\mathbf{V}_{i}$ are concatenated with all key and value vectors from all previous steps to form the complete key tensor and the value tensor. These key tensor and value tensor are cached and will be reused in the subsequent steps.
$\mathbf{K}$ =$=\begin{cases}\mathbf{K}_{i},&\text{if }i=1,\\ \mathrm{concat}\left(\mathbf{K}_{[1:i]},\mathbf{K}_{i}\right),&\text{otherwise},\end{cases}$ (88)
$\mathbf{V}$ =$=\begin{cases}\mathbf{V}_{i},&\text{if }i=1,\\ \mathrm{concat}\left(\mathbf{V}_{[1:i]},\mathbf{V}_{i}\right),&\text{otherwise},\end{cases}$
In this way, the attention computation at each step $i$ leverages the cached key-value pairs from previous steps can be written as
$\mathbf{z}_{i}=\mathrm{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V},$ (89)
significantly reducing redundant operations and improving inference efficiency in autoregressive decoding.
In autoregressive models, the use of causal attention masks ensures that caching is theoretically lossless. This is attributed to the unidirectional nature of attention, where each token attends only to its preceding tokens. Consequently, previously computed key and value representations remain valid and unchanged throughout the generation process. Similarly, in semi-autoregressive generation paradigms such as block-wise decoding , caching can be applied between different blocks without introducing approximation errors.
In contrast, dLLMs and dMLLMs employ full (bidirectional) attention mechanisms, wherein each token can attend to all other positions, regardless of their masking status. As a result, even tokens that have already been decoded and unmasked may have their key and value vectors influenced by updates to other tokens during subsequent diffusion iterations.
Despite this theoretical limitation, empirical studies have consistently validated the effectiveness of incorporating prefilling and KV cache techniques in dLLMs and dLLMs. These methods yield substantial acceleration in inference with only minimal degradation in model performance.
V.C.1 Prefilling
In multimodal large models, the inclusion of visual inputs significantly increases the number of prompt tokens, often exceeding the length of the response even without inference. Prefilling the prompt greatly enhances inference efficiency. The papers Dimple [8] and LaViDa [10] were among the first to apply prefilling techniques to dMLLMs, demonstrating the feasibility of this approach in multimodal large models.
Dimple‘s experimental results demonstrate that the use of prefilling incurs negligible performance degradation on the majority of vision-language benchmarks. This observation implies that current multimodal models process and encode visual information in a largely static, one-pass manner, with limited dependence on the subsequent generation of textual responses. Furthermore, the adoption of prefilling yields substantial improvements in inference efficiency, achieving speedups ranging from 2× to 7×. Notably, the magnitude of acceleration increases with larger batch sizes, indicating a positive correlation between speedup and GPU utilization.
V.C.2 KV-Cache
Due to the use of bidirectional attention, the cached KV pairs, attention outputs, and other values in dLLMs and dMLLMs are not static. The associated caching algorithms typically consist of three components: caching, reuse, and update.
@@ V.C.2 KV-Cache
  • dKV-Cache [25]. The core idea of dKV-Cache is to cache the key-value pairs when the tokens are unmasked and reuse the cached key-value pairs. Additionally, it introduces an interval hyperparameter to periodically update the cached key-value pairs.
@@
The general experimental observations are that caching in dLLMs leads to performance fluctuations, and in most cases, results in performance degradation. The severity of degeneration increases with larger update intervals. However, with smaller update intervals (e.g., 2 to 8), performance degradation is minimal and often comparable to that of models without caching [25, 26].
V.D Guidance Techniques
In dLLMs and dMLLMs, post-processing on the predicted logits or sampling probabilities is commonly referred to as guidance, following terminology from image diffusion models. Guidance methods are used to influence the generation process toward desired characteristics, such as improved diversity or controllability.
V.D.1 Classifier-Free Guidance
[19] proposes an unsupervised classifier-free guidance strategy for discrete diffusion generation. The method performs two forward predictions at each diffusion timestep $t$:
@@ V.D.1 Classifier-Free Guidance
V.D.2 Classifier Guidance
To improve controllability, for block diffusion model, [52] introduces a classifier guidance framework that integrates class-conditional preferences into the sampling process.

### v4

@@ V.E Addressing Training-Testing Discrepancy
V.F Reweighting Technique
@@
V.G Distillation through Dimensional Correlations
@@
$\mathcal{L}_{\text{distil}}(\theta)$ $=\mathbb{E}_{x_{\delta}\sim r_{\delta}}\Big{[}\mathrm{D_{KL}}\big{(}p^{\psi}_{0|\delta}(\cdot\mid x_{\delta})p^{\theta}_{0|\delta}(\cdot\mid x_{\delta})\big{)}\Big{]}$ (81)
@@
$\mathcal{L}_{\text{consis}}(\theta)={}\mathbb{E}_{x_{t}\sim r_{t}}\Big{[}\mathrm{D_{KL}}\bigl{(}(p^{\theta}_{s|u}\circ p^{\psi}_{u|t})(\cdot\mid x_{t})\|\,p^{\theta}_{s|t}(\cdot\mid x_{t})\bigr{)}\Big{]}$ (82)
@@
V.H Reinforcement Learning
Reinforcement learning has been extensively applied to discrete diffusion language models [11, 24, 64, 63, 62, 148, 74]. Reinforcement learning for diffusion language models largely inherits the paradigm from autoregressive models; however, it also presents several unique challenges, such as the estimation of likelihood.
V.H.1 Diffusion-based GRPO
UniGRPO in [11] adapts policy optimization to the diffusion setting by combining structured noising, likelihood approximation, and clipped policy gradients. Let $q$ and $\{o_{i}\}_{i=1}^{G}$ denote a query and a batch of responses, respectively. For each response $o_{i}$, UniGRPO samples a masking ratio $p_{i}\sim U[0,1]$ and create a perturbed version $\tilde{o}_{i,p}$ by masking tokens. The token-level likelihood and the sequence-level likelihood are approximated as
$\pi^{\prime}_{\theta}(o_{i,t}\mid q,\tilde{o}_{i,p},p_{i})$ $=\mathbb{E}_{p_{i}}\big{[}\mathbf{1}[\tilde{o}_{i,t,p}=\texttt{[M]}]\log p_{\theta}(o_{i,t,p}\mid q)\big{]},$ (83)
$\pi^{\prime}_{i}$ $=\frac{1}{M}\sum_{o_{i,t}\in M}\log p_{\theta}(o_{i,t}\mid q).$ (84)
Following GRPO, UniGRPO loss is an integration of the clipped surrogate rewards $R_{i,t}$ with the KL regularization:
$\mathcal{J}_{\text{UniGRPO}}(\theta)=\mathbb{E}\!\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}R_{i,t}-\beta D_{\mathrm{KL}}(\pi^{\prime}_{\theta}\|\pi^{\prime}_{\mathrm{ref}})\right].$ (85)
V.H.2 Variance-Reduced Preference Optimization
LLaDA 1.5 [24] introduces the Variance-Reduced Preference Optimization (VRPO) which replaces the intractable log-likelihoods in DPO with Evidence Lower Bounds (ELBOs)
$\mathcal{B}_{\pi}(y\mid x)$ $\triangleq\mathbb{E}_{t\sim\mathcal{U}[0,1]}\mathbb{E}_{y_{t}\sim q(y_{t}\mid t,y,x)}\ell_{\pi}(y_{t},t,y\mid x)$ (86)
$\leq\log\pi(y\mid x),$ (87)
where $\ell_{\pi}$ is the mask prediction loss. To reduce estimator variance, VRPO introduces (i) increased sampling budgets, (ii) full sampling budget over timesteps, and (iii) antithetic sampling that shares noise between $\pi_{\theta}$ and $\pi_{\text{ref}}$.
V.I Stepwise Decomposition Preference Optimization
A central challenge in reinforcement learning alignment for discrete diffusion models is how to propagate reward information through the entire denoising trajectory. Stepwise Decomposition Preference Optimization (SDPO) [62] reformulates trajectory alignment into a collection of tractable per-step alignment objectives.
The standard KL-regularized reward optimization objective for diffusion trajectories is:
$\max_{p_{\theta}}\;\mathbb{E}_{p_{\theta}(x_{0:T}\mid c)}\big{[}r(x_{0},c)\big{]}-\beta D_{\mathrm{KL}}\!\big{[}p_{\theta}(x_{0:T}\mid c)\;\|\;p_{\mathrm{ref}}(x_{0:T}\mid c)\big{]},$ (88)
where $c$ is the conditioning context, $r(x_{0},c)$ is a reward function on the clean sequence, and $p_{\mathrm{ref}}$ is the reference model. However, this requires sampling and scoring the entire trajectory, which is computationally prohibitive. Instead of optimizing the whole trajectory, SDPO aligns the per-step posterior:
$p_{\theta}(x_{0}\mid x_{t},c),$ (89)
which admits exact likelihood evaluation under masked diffusion models. The stepwise alignment objective is:
$\max_{p_{\theta}}\mathbb{E}_{p_{\theta}(x_{0}\mid x_{t},c)}[r(x_{0},c)]-\beta_{t}D_{\mathrm{KL}}\big{[}p_{\theta}(x_{0}\mid x_{t},c)||p_{\mathrm{ref}}(x_{0}\mid x_{t},c)\big{]},$ (90)
with step-dependent regularization $\beta_{t}=\beta/w(t)$. The work shows that this step-wise optimization is equivalent to a distribution matching problem, which can be used to further simplify the loss function.
V.J Weighted Policy Optimization
Weighted Policy Optimization (wd1) [63] reformulates the reinforcement learning objective as a weighted likelihood objective.
Under the reverse-KL–regularized policy optimization, the target solution $\pi^{\ast}$ has the closed form. wd1 minimizes the KL divergence $D_{\mathrm{KL}}(\pi^{\ast}\,\|\,\pi_{\theta})$, which reduces to a weighted negative log-likelihood:
$L_{\mathrm{wd1}}(\theta)=-\mathbb{E}_{q\sim D,\{o_{i}\}\sim\pi^{\text{ref-old}}}\left[\sum_{i=1}^{G}w(q,o_{i})\cdot\log\pi_{\theta}(o_{i}\mid q)\right].$ (91)
To address the issue where samples with small advantages receive disproportionately low weights, the weights are defined as:
$w(q,o_{i})$ $=-w^{+}(q,o_{i})+w^{-}(q,o_{i}),$ (92)
$w^{+}(q,o_{i})$ $=\frac{\exp(\psi A_{i})}{\sum_{j=1}^{G}\exp(\psi A_{j})},$ (93)
$w^{-}(q,o_{i})$ $=\frac{\exp(-\psi A_{i})}{\sum_{j=1}^{G}\exp(-\psi A_{j})},$ (94)
where $A_{i}$ is the centered reward and $\psi$ is a scaling factor.
V.J.1 Diffusion Chain of Lateral Thought
Diffusion Chain of Lateral Thought (DCoLT) [64] introduces a Unmask Policy Module (UPM), which is trained via reinforcement learning to control the decoding order. The UPM learns a ranking-based policy over masked tokens:
$h^{i}_{\theta,t}=\text{UPM}(x_{t-1},t,m_{i}),$ (95)
where $h^{i}_{\theta,t}$ is the predicted score for token $i$ at step $t$, and $m_{i}$ is its mask indicator. A Plackett–Luce distribution is then used to sample a top-$K$ unmasking set $U_{t}$. Once $U_{t}$ is selected, the model predicts token values for those positions using the standard diffusion langauge model. The UPM is implemented as a lightweight transformer block attached to diffusion langauge model.
VI Inference Techniques
VI.A Unmasking Techniques
@@
In dLLMs and dMLLMs, the model predicts all the response tokens at each step. However, only a subset of the masked tokens are selected to be unmasked at every iteration, while the remainder remain masked. This iterative unmasking process gradually reveals the complete output sequence. The core challenges in this procedure are: (1) determining which tokens to unmask at each iteration, and (2) deciding how many tokens should be unmasked at each iteration. Figure 5 provides a detailed illustration of the unmasking strategies. In this section, we discuss each category in detail and describe the specific unmasking strategies proposed in each work.
VI.A.1 Discrete-Time Unmasking
@@
Let $p^{(i)}\in\mathbb{R}^{K}$ be the predicted probability distribution over the vocabulary for the $i$-th token, where $K$ is the vocabulary size. The following are some commonly used metrics.
  • Maximum Probability (Confidence) [7, 65]:
    $c^{(i)}=\max(p^{(i)}),$ (96)
@@
  • Margin [7, 65]:
    $c^{(i)}=p_{\text{top1}}^{(i)}-p_{\text{top2}}^{(i)},$ (97)
    where $p_{\text{top1}}^{(i)}$ and $p_{\text{top2}}^{(i)}$ are the first and second highest probabilities, respectively. This measures how dominant the top prediction is.
  • Negative Entropy [7, 65]:
    $c^{(i)}=-\sum_{k=1}^{K}p_{k}^{(i)}\log(p_{i}+\epsilon),$ (98)
@@
  • In EB-Sampler [66], instead of performing filtering at the token level, the metric is defined over a set of tokens, aiming to directly decide on unmasking a subset $U$ of tokens.
    $c^{(U)}=\sum_{l\in U}H(p^{(l)})-\max_{l\in U}H(p^{(l)})$ (99)
    with $H(p^{(l)})$ is the entropy of $p^{(l)}$.
  • In PC-Sampler. [67], the metric is defined as
    $c^{(i)}=w^{(i)}\cdot p^{(i)}_{x^{(i)}}\cdot[-\log p_{\mathcal{D}}(x^{(i)})],\quad w^{(i)}=e^{-\lambda\cdot i},$ (100)
    where $x^{(i)}$ is the predicted token at position $i$. In this metric, $w^{(i)}$ controls the degree of left-to-right decoding by applying an exponentially decaying weight based on token positions. The parameter $\lambda$ serves as a hyperparameter. $p^{(i)}_{x^{(i)}}$ denotes the model’s predicted probability for the current token, reflecting its confidence. $p_{\mathcal{D}}$ represents the token frequency distribution estimated from a publicly available corpus $p_{\mathcal{D}}$, used to downweight trivial words such as “the” and “a”.
@@
  • Block-wise Unmasking: Block-wise semi-autoregressive decoding strategy [6] divides the full response into multiple blocks, similar to block diffusion. During each forward pass, predictions are generated for all blocks simultaneously. However, the unmasking of tokens follows a left-to-right, block-by-block order. Tokens in the next block are only allowed to be unmasked once all tokens in the previous block have been unmasked.
  • Dilated Unmasking Schedule (DUS) [68]: DUS considers the relative distances between tokens during decoding and requires the distances among decoded tokens to decrease gradually from large to small. Specifically, let $\mathcal{U}_{t}$ denote the selected tokens at time $t$, DUS selects equally spaced tokens as following:
    $\mathcal{U}_{0}$ $=\varnothing,$ (102)
    $\mathcal{P}_{t}$ $=\{k\mid(k-1)\bmod s_{t}=0\},$ (103)
    $\mathcal{U}_{t}$ $=\mathcal{U}_{t-1}\cup\mathcal{P}_{t},$ (104)
    where $s_{t}$ is a dilation coefficient that decreases over time.
@@
VI.A.2 Continuous-Time Unmasking (Flow Matching)
@@
VI.B Remasking Techniques
For masked discrete diffusion model, once a token is unmasked, it remains unchanged in subsequent steps. This static behavior limits the model’s capacity to revise or refine earlier predictions. To address this, the remasking technique reintroduces masked tokens at previously unmasked positions, enabling iterative refinement of generated outputs.
VI.B.1 Remasking in General Masked Diffusion Models.
@@ VI.B Remasking Techniques
VI.B.2 Wide-In Narrow-Out
Under the block-wise decoding setting, [69] introduces an alternative verification-based remasking approach. Specifically, A appends an additional set of shadow tokens after the current response, with the number of shadow tokens equal to the current block size and each shadow token corresponding to one token in the block. By adjusting the attention mask, the original token sequence is prevented from attending to the shadow tokens, while each shadow token can attend to all tokens except its corresponding token in the current block. Consequently, the shadow tokens verify the decoded tokens without interfering with the decoding process of the original response sequence. If some decoded tokens are found to have low confidence, they are replaced with mask tokens.
VI.B.3 Remasking under Discrete Flow Matching.
@@ VI.B.2 Wide-In Narrow-Out
VI.C Prefilling and Caching Techniques
@@
Prefilling and Key-Value Cache (KV-Cache) are standard inference acceleration techniques widely adopted in autoregressive language models. Intuitively, Prefilling and KV-Cache avoids redundant computation by storing the key and value representations from previous decoding steps, enabling the model to reuse them instead of recalculating at each new time step. In autoregressive models, the use of causal attention masks ensures that caching is theoretically lossless. In contrast, dLLMs and dMLLMs employ full (bidirectional) attention mechanisms, wherein each token can attend to all other positions. As a result, even tokens that have already been decoded and unmasked may have their key and value vectors influenced by updates to other tokens during subsequent diffusion iterations. Thus, caching in dLLM and dMLLM are not theoretically lossless.
For dLLMs, dKV-Cache [29] and dLLM-Cache [30] develops the naive KV-Cache techniques. Their observations are that, with small update intervals (e.g., 2 to 8), caching in dLLMs leads to minimal performance degradation and achieves 1̃0x speed-up. For dMLLMs, Dimple [8] and LaViDa [10] empirically verify that the use of prefilling incurs negligible performance degradation on the majority of vision-language benchmarks and provides a 2× to 7× speed-up.
Following are some representative designs for KV-Cache in dLLM.
@@
  • dKV-Cache [29]. The core idea of dKV-Cache is to cache the key-value pairs when the tokens are unmasked and reuse the cached key-value pairs with an interval hyperparameter to periodically update the cached key-value pairs.
@@
VI.D Guidance Techniques
In dLLMs and dMLLMs, post-processing on the predicted logits or sampling probabilities is commonly referred to as guidance, following terminology from image diffusion models.
VI.D.1 Classifier-Free Guidance
[149, 22] introduce the unsupervised classifier-free guidance strategy for discrete diffusion generation. The method performs two forward predictions at each diffusion timestep $t$:
@@ VI.D.1 Classifier-Free Guidance
In the continuous-time setting, [150] provides the formulation of classifier-free guidance applied to the reverse transition matrix
$\hat{R}_{t}(x,x^{\prime})=R_{t}(x^{\prime},x)\left(\frac{p_{t}(x)}{p_{t}(x^{\prime})}\right)^{w}\left(\frac{q_{t}(x)}{q_{t}(x^{\prime})}\right)^{1-w},$ (108)
where distribution $q_{t}$ is the reference distribution. [151] further demonstrate that applying strong guidance during early decoding can severely degrade sample quality. This degradation arises because high guidance accelerates the unmasking rate, leading to overly confident and premature token predictions. To mitigate this issue, the authors suggests applying a column-wise normalization of the guided transition matrix.
@@
VI.D.2 Classifier Guidance
To improve controllability, for block diffusion model, [149, 71] introduce the classifier guidance framework that integrates class-conditional preferences into the sampling process.